# Implicit Bias-Like Patterns in Reasoning Models

### v2

@@ Abstract
Implicit bias refers to automatic mental processes that shape perceptions, judgments, and behaviors. Previous research on “implicit bias" in LLMs focused primarily on outputs rather than the processes underlying the outputs. We present the Reasoning Model Implicit Association Test (RM-IAT) to study implicit bias-like processing in reasoning models, which are LLMs using step-by-step reasoning for complex tasks. Using RM-IAT, we find o3-mini and DeepSeek R1 require more tokens when processing association-incompatible information, mirroring human implicit bias patterns. Conversely, Claude 3.7 Sonnet displays reversed patterns for race and gender tests, requiring more tokens for association-compatible information. This reversal appears linked to differences in safety mechanism activation, increasing deliberation in sensitive contexts. These findings suggest AI systems can exhibit processing patterns analogous to both human implicit bias and bias correction mechanisms.
Introduction
@@
Implicit bias refers to automatic mental processes that shape perceptions, judgments, and behaviors based on social categories such as race, gender, or age Greenwald and Lai (2020); Payne and Gawronski (2010). Implicit biases often operate rapidly and with high efficiency, requiring minimal cognitive resources while influencing judgments through the automatic activation of stored information about social groups Melnikoff and Bargh (2018); Bargh and Williams (2006); Fazio et al. (1986). This efficiency in processing allows implicit biases to operate even under conditions of limited attention or cognitive load. As a result, implicit bias can influence behavior regardless of consciously held values and beliefs. Research demonstrates that implicit bias significantly relates to real-world outcomes, with researchers describing a potential role of implicit bias in domains such as employment Agerström and Rooth (2011), healthcare FitzGerald and Hurst (2017), and criminal justice Spencer et al. (2016).
Implicit Association Test (IAT)
@@
To measure these automatic evaluations in humans, social psychologists developed the Implicit Association Test (IAT; Greenwald et al. (1998)). The IAT has become the most popular measure for assessing implicit biases and has cited and used in thousands of papers Greenwald and Lai (2020). During the test, participants are instructed to rapidly pair group category stimuli with attributes. For example, in the Race IAT, participants are asked to press one key for White faces and pleasant words and another for Black faces and unpleasant words (i.e., association-compatible pairings). After many trials of pairing stimuli to those categories, the pairings are switched. In the next set of trials, Black faces and pleasant words share a key, while White faces and unpleasant words share the other key (i.e., association-incompatible pairings). The difference in response times between these sets of trials informs understanding about how concepts are linked together in memory. Typically, people show faster responses for association-compatible pairings than incompatible pairings, indicating the presence of automatically activated associations between social groups and stereotypical attributes.
Implicit bias in language models
@@
With recent advances in language models, researchers have explored whether language models exhibit biases like humans do. Past work has focused on examining bias in how language models reproduce societal stereotypes in their generated content Abid et al. (2021); Lucy and Bamman (2021). In light of these findings, more recent models undergo extensive post-training steps such as instruction fine-tuning and supervised learning to ensure that models align with human values Ziegler et al. (2020); Ouyang et al. (2022). As a result, the more recent models are less likely to express bias in generated content.
@@
However, there remain concerns about implicit bias in language models. Zhao et al. (2024) had GPT-3.5 complete templates by filling in social group pairs (e.g., “X are nurses as Y are surgeons") and then evaluating those completions as “right" or “wrong." Models tended to generate stereotypical completions (e.g., “Women are nurses as Men are surgeons") while simultaneously labeling them as “wrong." Bai et al. (2025) administered a modified version of the Implicit Association Test, asking LLMs to pick words signaling social group identities (e.g., Julia and Ben) next to a list of attribute words (e.g., home, work). Then, they calculated the proportion of association-compatible pairings. The authors found that proprietary LLMs, despite being trained to align with human values and avoid expressing biases, still showed a stronger tendency to create association-compatible rather than incompatible pairings.
@@
With advanced post-training techniques making biases harder to detect in model outputs, attention has shifted to more subtle forms of biases in how LLMs operate. These patterns, while not detectable in conventional evaluations, may systematically influence model behavior in consequential ways. Bai et al. (2025) demonstrated this by showing a correlation between LLMs’ responses in a modified Implicit Association Test and models’ tendency to exhibit bias in decision-making contexts. Given the growing deployment of language models in decision-making contexts, understanding and addressing these hard-to-detect patterns is crucial for preventing discriminatory outcomes, particularly as models become more proficient at avoiding blatant forms of bias.
Reasoning models and reasoning tokens
@@
Reasoning models represent a breakthrough in language model capabilities. Through reinforcement learning, reasoning models have been trained to generate intermediate reasoning steps, producing a sequence of “reasoning tokens" that represent their thought process. These reasoning tokens are step-by-step articulations of the model’s problem-solving approach before it arrives at a final answer. For example, when asked “What is 23 $\times$ 48?", a reasoning model might generate tokens like: “Let me break this down: (20 + 3) × 48 = 20 $\times$ 48 + 3 $\times$ 48 = 960 + 144 = 1104." This reasoning approach has been shown to dramatically improve model performance on complex tasks such as coding, commonsense and arithmetic reasoning Wei et al. (2023).
@@
Reasoning tokens provide a unique window for researchers to understand how models process information, paralleling how response latencies are used to quantify automatic evaluations in IATs. Similar to how increased response times in humans indicate greater cognitive effort and deliberation when processing association-incompatible information, a higher reasoning token count suggests increased computational processing when the model encounters associations that contradict previously observed patterns. This metric enables assessment of automatic evaluations, offering a closer parallel to the IAT.
This work
@@
Previous research purporting to measure implicit bias patterns in language models has primarily examined model outputs and word associations Bai et al. (2025); Zhao et al. (2024). However, these approaches capture the outcomes of how a model processes information rather than the actual processing of information itself. Such associations could simply reflect biases present in training data rather than computational patterns analogous to human implicit cognition. They also cannot capture the speed and efficiency at which information is processed, which is a key aspect of implicit bias in research on humans Bargh (1994); Greenwald and Lai (2020). In this work, we propose a novel approach that examines the degree of automaticity or deliberation in reasoning models by measuring how much models expend computational effort through reasoning token usage. This framework parallels how human implicit bias is studied through response latencies in the IAT. In the IAT, processing efficiency decreases when handling association-incompatible information, requiring greater effort that result in delays in response time. By analyzing computational processing patterns (i.e., how a model “thinks") rather than outputs (i.e, what a model “does"), we more closely capture phenomena analogous to implicit bias in language models.
@@
Inspired by work on human implicit bias, we adapted the Implicit Association Test into the RM-IAT (Reasoning Model Implicit Association Test) to gauge how much deliberation a reasoning model invests when processing association-compatible versus incompatible information. In the RM-IAT, the reasoning model is first shown the complete list of words that represent the two group categories (e.g., men and women) and the two attribute categories (e.g., career and family). Once the stimuli has been introduced, the model receives a prompt asking it to assign an individual group word (e.g., “John") to one of the attribute categories. The experiment consists of two conditions. In the association-compatible condition, the prompt instructs the model to categorize group words in a way that aligns with established associations (e.g., men and career / women and family). In the association-incompatible condition, the model is instructed to categorize group words in a way that conflicts with established associations (e.g., women and career / men and family). For every categorization, we record the model’s reasoning token count–the number of tokens generated in its reasoning step before answering. Then we compare the average token counts across the two conditions just as the human IAT compares response latencies. Higher token counts mirror longer response times in IATs, indicating greater computational effort on a specific condition (see Figure 1).
@@
We administered the RM-IAT using three state-of-the-art reasoning models–o3-mini OpenAI (2025), DeepSeek R1 DeepSeek-AI (2025), and Claude 3.7 Sonnet Anthropic (2025). Both o3-mini and DeepSeek R1 exhibited clear patterns, requiring significantly more reasoning tokens to process association-incompatible pairings than association-compatible pairings in most RM-IATs (9 out of 10 for o3-mini and 8 out of 10 for DeepSeek R1). Claude 3.7 Sonnet displayed more complex behavior, showing reversed patterns in RM-IATs related to race and gender while exhibiting consistent association patterns in 3 RM-IATs that were about less socially sensitive topics.
Methods
@@
Reasoning Model Selection and Reasoning Tokens
We used three state-of-the-art reasoning models for data collection: o3-mini OpenAI (2025), DeepSeek R1 DeepSeek-AI (2025), and Claude 3.7 Sonnet with extended thinking Anthropic (2025). We made 12,920 API calls for each model, collecting both the model’s final categorization response and the number of tokens used in its reasoning chain from each API call. The API provides reasoning token counts, offering a quantitative measure of computational processing required for each categorization. [^1] For the o3-mini model specifically, all 12,920 OpenAI API calls produced token counts that were multiples of 64, suggesting either hardware constraints or model-specific instructions for reasoning (i.e., the model may be configured to process reasoning in 64-token increments). This approximation introduces imprecision in our measurements for o3-mini, potentially obscuring the true magnitude of differences between conditions. For Claude 3.7 Sonnet and DeepSeek R1 whose reasoning tokens were accessible via their API, the tokens were collected too. [^2] Claude 3.7 Sonnet and DeepSeek R1 do not separate reasoning tokens from final output tokens. Since the final output was limited to the categorization task, we simply used the output_tokens as our measurement of reasoning token count. Default parameters were used. [^3] The default value of the reasoning_effort parameter–which determines the number of tokens OpenAI reasoning models use during reasoning–is “medium.” Claude 3.7 Sonnet requires that the maximum number of tokens that the model can use for the task be set a priori, which we set to 5,020, far exceeding the maximum number of reasoning tokens that was used by o3-mini–2,304.
Prompting the Implicit Association Test
@@
The traditional IAT consists of seven blocks Greenwald et al. (2009, 2003). The first two blocks participants are familiarized with the task by classifying labels or images used to represent two group categories (e.g., names of men and women) and two attribute categories (e.g., career and family). In the first set of combined blocks (Blocks 3-4), participants respond to association-compatible pairings by pressing the same key for instruments/pleasant and weapons/unpleasant. After a practice block where only the two categories are presented and switched (Block 5), the second set of combined blocks (Blocks 6-7) presents association-incompatible pairings (e.g., same key for instruments/unpleasant and weapons/pleasant). Blocks 3-4 and 6-7 are usually counterbalanced to control for possible order effects. Measures of implicit bias are computed by comparing mean reaction times on the association-compatible blocks against mean reaction times on the association-incompatible blocks.
@@
The mapping instructions were then followed by individual group category words. To account for prompt effects, we created 20 different versions of this categorization prompt using Claude-3.5 Anthropic (2024). Each version maintained the same basic task–asking the model to assign a word to one of two attribute categories–while varying the phrasing. The first ten prompt variations can be found in Table S2 of the Supplementary Materials.
@@
Category and Stimulus Selection
@@
Caliskan et al. (2017) tested for human-like biases in word embedding models, which represent words as numeric vectors encoding semantic meaning. They examined past work from the social psychology literature, most using the IAT, and extracted word stimuli from these studies to test for 10 different associations in word embeddings. However, some low-frequency words were removed from their analyses as the model didn’t return representations for those words. Since reasoning models don’t share this limitation, we used all original word stimuli, found in Table S1 of the Supplementary Materials. In each of the original IATs and Caliskan et al.’s study, they found biases wherein pairings between first target and the first attribute (e.g., Flowers + Pleasant) and the second target and the second attribute (e.g., Insects + Unpleasant) were more compatible than the reverse (e.g., Flowers + Unpleasant / Insects + Pleasant).
  • Flowers/Insects + Pleasant/Unpleasant from Greenwald et al. (1998)
  • Instruments/Weapons + Pleasant/Unpleasant from Greenwald et al. (1998)
  • European/African Americans + Pleasant/Unpleasant (1) from Greenwald et al. (1998)
@@
  • European/African Americans + Pleasant/Unpleasant (2) from Bertrand and Mullainathan (2004) [^4] This was not an IAT study. We used the list of names from their experiment rather than those in Greenwald et al. (1998).
  • European/African Americans + Pleasant/Unpleasant (3) from Nosek et al. (2002a)
  • Men/Women + Career/Family from Nosek et al. (2002a)
  • Men/Women + Mathematics/Arts from Nosek et al. (2002b)
  • Men/Women + Science/Arts from Nosek et al. (2002a)
@@
  • Mental/Physical Diseases + Temporary/Permanent from Monteith and Pettit (2011)
  • Young/Old People + Pleasant/Unpleasant from Nosek et al. (2002a)
Comparison of the Number of Reasoning Tokens
[…]

### v3

@@ Abstract
Implicit biases refer to automatic mental processes that shape perceptions, judgments, and behaviors. Previous research on “implicit bias” in LLMs focused primarily on outputs rather than the processes underlying the outputs. We present the Reasoning Model Implicit Association Test (RM-IAT) to study implicit bias-like processing in reasoning models, which are LLMs that use step-by-step reasoning for complex tasks. Using RM-IAT, we find that reasoning models like o3-mini, DeepSeek-R1, gpt-oss-20b, and Qwen-3 8B consistently expend more reasoning tokens on association-incompatible tasks than association-compatible tasks, suggesting greater computational effort when processing counter-stereotypical information. In contrast, Claude 3.7 Sonnet exhibited reversed or inconsistent patterns, likely due to embedded safety mechanisms that flagged or rejected socially sensitive associations. These divergent behaviors highlight important differences in how alignment and safety processes shape model reasoning. As reasoning models become increasingly integrated into real-world decision-making, understanding their implicit bias-like patterns and how alignment methods influence them is crucial for ensuring fair and trustworthy AI systems.
1 Introduction
@@
Implicit biases refer to automatic mental processes that shape perceptions, judgments, and behaviors based on social categories such as race, gender, or age (Greenwald and Lai, 2020; Payne and Gawronski, 2010). Implicit biases often operate rapidly and with high efficiency, requiring minimal cognitive resources while influencing judgments through the automatic mental activation of information about social groups (Melnikoff and Bargh, 2018; Bargh and Williams, 2006; Fazio et al., 1986). This efficiency in processing means implicit biases operate even when attention is limited and deliberation is non-existent. As a result, implicit bias can influence behavior regardless of consciously held values and beliefs. Research demonstrates that implicit bias significantly relates to real-world outcomes, with researchers describing a potential role of implicit bias in domains such as employment (Agerström and Rooth, 2011), healthcare (FitzGerald and Hurst, 2017), and criminal justice (Spencer et al., 2016).
1.1 Implicit Association Test (IAT)
@@
To measure these automatic evaluations in humans, social psychologists developed the Implicit Association Test (IAT; Greenwald et al. (1998)). The IAT has become the most popular measure for assessing implicit biases and has cited and used in thousands of papers (Greenwald and Lai, 2020). During the test, participants are instructed to rapidly pair group category stimuli with attributes. For example, in the Race IAT, participants are asked to press one key for White faces and pleasant words and another for Black faces and unpleasant words (i.e., association-compatible pairings). After many trials of pairing stimuli to those categories, the pairings are switched. In the next set of trials, Black faces and pleasant words share a key, while White faces and unpleasant words share the other key (i.e., association-incompatible pairings). The difference in response times between these sets of trials informs understanding about how concepts are linked together in memory. People show faster responses for association-compatible pairings than incompatible pairings, indicating the presence of automatically activated associations between social groups and stereotypical attributes.
The IAT is best positioned to capture the efficiency of mental processing (Gawronski, 2024; Goedderz et al., 2024; Lai and Wilson, 2021; Morris and Kurdi, 2023). According to an influential model called the Iterative Reprocessing Model (Cunningham and Zelazo, 2007), initial evaluations of stimuli rely on readily accessible information and require minimal mental effort. As the IAT requires rapid responding, the IAT is especially sensitive to these efficient processes. The Iterative Reprocessing Model also theorizes that as more elaborate and complex information becomes available, people will reinterpret stimuli in a more deliberate manner. Measures that allow for less mentally efficient responding, like self-report surveys, typically capture these later stages of processing where more complex information is considered.
1.2 Implicit bias in language models
@@
With recent advances in language models, researchers have explored whether language models exhibit biases like humans do. Past work has focused on examining bias in how language models reproduce societal stereotypes in their generated content (Abid et al., 2021; Lucy and Bamman, 2021). In light of these findings, more recent models undergo extensive post-training steps such as instruction fine-tuning and supervised learning to ensure that models align with human values (Ziegler et al., 2020; Ouyang et al., 2022). As a result, the more recent models are less likely to express bias in generated content.
@@
However, there remain concerns about implicit bias in language models. (Zhao et al., 2024) had GPT-3.5 complete templates by filling in social group pairs (e.g., “X are nurses as Y are surgeons") and then evaluating those completions as “right" or “wrong." Models tended to generate stereotypical completions (e.g., “Women are nurses as Men are surgeons") while simultaneously labeling them as “wrong." (Bai et al., 2025) administered a modified version of the Implicit Association Test, asking LLMs to pick words signaling social group identities (e.g., Julia and Ben) next to a list of attribute words (e.g., home, work). Then, they calculated the proportion of association-compatible pairings. The authors found that proprietary LLMs, despite being trained to align with human values and avoid expressing biases, still showed a stronger tendency to create association-compatible rather than incompatible pairings.
@@
With advanced post-training techniques making biases harder to detect in model outputs, attention has shifted to more subtle forms of biases in how LLMs operate. These patterns, while not detectable in conventional evaluations, may systematically influence model behavior in consequential ways. (Bai et al., 2025) demonstrated this by showing a correlation between LLMs’ responses in a modified Implicit Association Test and models’ tendency to exhibit bias in decision-making contexts. Given the growing deployment of language models in decision-making contexts, understanding and addressing these hard-to-detect patterns is crucial for preventing discriminatory outcomes, particularly as models become more proficient at avoiding blatant forms of bias.
1.3 Reasoning models and reasoning tokens
@@
Reasoning models represent a breakthrough in language model capabilities. Through reinforcement learning, reasoning models have been trained to generate intermediate reasoning steps, producing a sequence of “reasoning tokens" that represent their thought process. These reasoning tokens are step-by-step articulations of the model’s problem-solving approach before it arrives at a final answer. For example, when asked “What is 23 $\times$ 48?", a reasoning model might generate tokens like: “Let me break this down: (20 + 3) × 48 = 20 $\times$ 48 + 3 $\times$ 48 = 960 + 144 = 1104." This reasoning approach has been shown to dramatically improve model performance on complex tasks such as coding, commonsense and arithmetic reasoning (Wei et al., 2023).
@@
Reasoning tokens provide a unique window for researchers to understand how models process information, paralleling how response latencies are used to quantify automatic evaluations in IATs. Similar to how increased response times in humans indicate greater cognitive effort and deliberation when processing association-incompatible information, a higher reasoning token count suggests increased computational processing when the model encounters associations that contradict previously observed patterns. These reasoning models perform computations for every token they generate, with each reasoning token corresponding to a discrete series of computational operations. This token-level processing architecture means that reasoning token counts directly quantify the computational resources allocated to the problem-solving process, enabling assessment of processing efficiency and offering a closer parallel to the IAT’s assessment of efficient mental processing.
1.4 This work
@@
Previous research purporting to measure implicit bias patterns in language models has primarily examined model outputs and word associations (Bai et al., 2025; Zhao et al., 2024). However, these approaches capture the outcomes of how a model processes information rather than the actual processing of information itself. Such associations could simply reflect biases present in training data rather than computational patterns analogous to human implicit cognition. They also cannot capture the speed and efficiency at which information is processed, which is a key aspect of implicit bias in research on humans (Bargh, 1994; Greenwald and Lai, 2020). In this work, we propose a novel approach that examines the degree of automaticity or deliberation in reasoning models by measuring how much models expend computational effort through reasoning token usage. This framework parallels how human implicit bias is studied through response latencies in the IAT. In the IAT, processing efficiency decreases when handling association-incompatible information, requiring greater effort that result in delays in response time. By analyzing computational processing patterns (i.e., how a model “thinks") rather than outputs (i.e, what a model “does"), we more closely capture phenomena analogous to implicit bias in language models.
@@
Inspired by work on human implicit bias, we adapted the Implicit Association Test into the RM-IAT (Reasoning Model Implicit Association Test) to gauge how much deliberation a reasoning model invests when processing association-compatible versus incompatible information. In the RM-IAT, the reasoning model is first shown the complete list of words that represent the two group categories (e.g., men and women) and the two attribute categories (e.g., career and family). Once the stimuli has been introduced, the model receives a prompt asking it to assign an individual group word (e.g., “John") to one of the attribute categories. The experiment consists of two conditions. In the association-compatible condition, the prompt instructs the model to categorize group words in a way that aligns with established associations (e.g., men and career / women and family). In the association-incompatible condition, the model is instructed to categorize group words in a way that conflicts with established associations (e.g., women and career / men and family). For every categorization, we record the model’s reasoning token count–the number of tokens generated in its reasoning step before answering. Then we compare the average token counts across the two conditions just as the human IAT compares response latencies. Higher token counts mirror longer response times in IATs, indicating greater computational effort on a specific condition (see Figure 1). Like the IAT, we theorized the RM-IAT was well-positioned to capture implicit biases in terms of efficiency of processing.
@@
We administered the RM-IAT using five state-of-the-art reasoning models–o3-mini (OpenAI, 2025), DeepSeek-R1 (DeepSeek-AI, 2025), Claude 3.7 Sonnet (Anthropic, 2025), gpt-oss-20b (OpenAI et al., 2025), and Qwen-3 8B (Yang et al., 2025). o3-mini, DeepSeek-R1, gpt-oss-20b, and Qwen-3 8B exhibited clear patterns, requiring significantly more reasoning tokens to process association-incompatible pairings than association-compatible pairings in most RM-IATs. Claude 3.7 Sonnet displayed more complex behavior, showing reversed patterns in RM-IATs related to race and gender while exhibiting consistent association patterns in 3 RM-IATs that were about less socially sensitive topics.
2 Methods
@@
2.1 Reasoning Model Selection and Reasoning Tokens
We used three proprietary–o3-mini (OpenAI, 2025), DeepSeek-R1 (DeepSeek-AI, 2025), and Claude 3.7 Sonnet with extended thinking (Anthropic, 2025)–and two open-source–gpt-oss-20b (OpenAI et al., 2025) and Qwen-3 8B (Yang et al., 2025)–reasoning models for data collection. We made 12,920 API calls for each proprietary model, collecting both the model’s final categorization response and the number of tokens used in its reasoning chain from each API call. For the open-source models, we conducted additional data collection by running the models locally.
Each reasoning model processes reasoning and output tokens differently, but all provide separable measurements. For o3-mini, the OpenAI API provides a detailed breakdown of tokens used, including the number of reasoning tokens.111see https://platform.openai.com/docs/guides/reasoning. All 12,920 o3-mini API calls produced reasoning token counts that were multiples of 64, suggesting either hardware constraints or model-specific processing configurations. This introduces measurement error where actual token counts are rounded up to the next multiple of 64. For example, true counts of 24 and 63 would both be reported as 64, understating actual differences, while counts of 64 and 65 would be reported as 64 and 128 respectively, overstating the difference. However, this measurement error does not substantially impact our analysis given the large effect sizes observed and the fact that the rounding affects both experimental conditions equally. For Claude 3.7 Sonnet, the API provides the number of output tokens but not a separate count of reasoning tokens. However, since final responses were only 1-2 tokens long and both experimental conditions generated similar output lengths, this did not affect our measurements. Similarly, DeepSeek-R1 provides completion token counts that include both reasoning and final response tokens, but the minimal final response length makes this distinction negligible for our analysis. For the open-source models, reasoning tokens are clearly demarcated: Qwen-3 8B generates reasoning tokens within <think> and </think> tags, while gpt-oss-20b places reasoning tokens between the analysis tag and the assistantfinal tag.
We used default parameters for all models. The default value of the reasoning_effort parameter for o3-mini is “medium." Claude 3.7 Sonnet requires setting the maximum number of tokens a priori, which we set to 5,020, far exceeding the maximum number of reasoning tokens used by o3-mini (2,304).
2.2 Prompting the Implicit Association Test
@@
The traditional IAT consists of seven blocks (Greenwald et al., 2009, 2003). The first two blocks participants are familiarized with the task by classifying labels or images used to represent two group categories (e.g., names of men and women) and two attribute categories (e.g., career and family). In the first set of combined blocks (Blocks 3-4), participants respond to association-compatible pairings by pressing the same key for instruments/pleasant and weapons/unpleasant. After a practice block where only the two categories are presented and switched (Block 5), the second set of combined blocks (Blocks 6-7) presents association-incompatible pairings (e.g., same key for instruments/unpleasant and weapons/pleasant). Blocks 3-4 and 6-7 are usually counterbalanced to control for possible order effects. Measures of implicit bias are computed by comparing mean reaction times on the association-compatible blocks against mean reaction times on the association-incompatible blocks.
@@
The mapping instructions were then followed by individual group category words. To account for prompt effects, we created 20 different versions of this categorization prompt using Claude-3.5 (Anthropic, 2024). Each version maintained the same basic task–asking the model to assign a word to one of two attribute categories–while varying the phrasing. The first ten prompt variations can be found in Table S2 of the Supplementary Materials.
@@
2.3 Category and Stimulus Selection
@@
(Caliskan et al., 2017) tested for human-like biases in word embedding models, which represent words as numeric vectors encoding semantic meaning. They examined past work from the social psychology literature, most using the IAT, and extracted word stimuli from these studies to test for 10 different associations in word embeddings. However, some low-frequency words were removed from their analyses as the model didn’t return representations for those words. Since reasoning models don’t share this limitation, we used all original word stimuli, found in Table S1 of the Supplementary Materials. In each of the original IATs and Caliskan et al.’s study, they found biases wherein pairings between first target and the first attribute (e.g., Flowers + Pleasant) and the second target and the second attribute (e.g., Insects + Unpleasant) were more compatible than the reverse (e.g., Flowers + Unpleasant / Insects + Pleasant).
  • Flowers/Insects + Pleasant/Unpleasant from (Greenwald et al., 1998)
  • Instruments/Weapons + Pleasant/Unpleasant from (Greenwald et al., 1998)
  • European/African Americans + Pleasant/Unpleasant (1) from (Greenwald et al., 1998)
@@
  • European/African Americans + Pleasant/Unpleasant (2) from (Bertrand and Mullainathan, 2004) [^2] This was not an IAT study. We used the list of names from their experiment rather than those in (Greenwald et al., 1998).
  • European/African Americans + Pleasant/Unpleasant (3) from (Nosek et al., 2002a)
  • Men/Women + Career/Family from (Nosek et al., 2002a)
  • Men/Women + Mathematics/Arts from (Nosek et al., 2002b)
  • Men/Women + Science/Arts from (Nosek et al., 2002a)
@@
  • Mental/Physical Diseases + Temporary/Permanent from (Monteith and Pettit, 2011)
  • Young/Old People + Pleasant/Unpleasant from (Nosek et al., 2002a)
2.4 Comparison of the Number of Reasoning Tokens